# RNA-FrameFlow: Flow Matching for de novo 3D RNA Backbone Design

### v1

@@ 4.1 Global evaluation of generated RNA backbones
We begin by analyzing RNA-FrameFlow’s samples using the aforementioned evaluation metrics. For validity, we report percentage of samples with scTM $\geq 0.45$; for diversity, we report the ratio of unique structural clusters to total valid samples; and for novelty, we report the highest average pdbTM to a match from the PDB. For each sequence length between 40 and 150, at intervals of 10, we generate 50 backbones. Table 1 reports these metrics across different variants for the number of denoising steps $N_{T}$. We compare our model to protein-RNA-DNA complex co-design model MMDiff (Morehead et al., 2023), which is a diffusion model. As the original version of MMDiff was trained on shorted RNA sequences, we retrain it on our training set. Additionally, we inverse-folded MMDiff’s backbones using gRNAde.
We identify $N_{T}=50$ as the best-performing model that balances validity, diversity, and novelty; furthermore, it takes 4.74 seconds (averaged over 5 runs) to sample a backbone of length 100, as opposed to 27.3 seconds for MMDiff with 100 diffusion steps. We note that increasing $N_{T}$ does not improve validity despite allowing the model to perform more updates to atomic coordinate placements. Our model also outperforms MMDiff. On manual inspection, samples from MMDiff had significant chain breaks and disconnected floating strands; see Appendix D.1.
Table 1: Unconditional RNA backbone generation. We evaluate the performance of RNA-FrameFlow for multiple values for denoising steps $N_{T}$. The best-performing model uses $N_{T}=50$ steps, taking 4.74s to sample a backbone of length 100. Diffusion-based MMDiff generated no valid backbones and took 5×5$\times$5 × longer to sample.
@@ 4.2 Local evaluation with structural measurements
For our best-performing model with number of timesteps $N_{T}=50$, we plot histograms of bond distance, bond angles, and dihedral angles in Figure 4. We include the Earth Mover’s distance (EMD) between measurements from the training and generated distributions as an indicator of local realism (using 30 bins for each quantity). An ideal generative model will score an EMD close to 0.0 (i.e. consistent with the training set comprising naturally occurring RNA). In Table 3, we observe EMD values from our best-performing model’s backbones being significantly closer to 0.0 compared to MMDiff and random Gaussian all-atom point clouds (akin to an untrained model), which serve as sanity checks. We include histograms for MMDiff in Appendix D.1.
We also show RNA Ramachandran angle plots for generated samples and the training distribution in Figure 4. Keating et al. (2011) introduced $\eta-\theta$ plots, similar to Ramachandran angle plots for proteins, that track the separate dihedral angles formed by $\{C4^{\prime}_{i},P_{i+1},C4^{\prime}_{i+1},P_{i+2}\}$ and $\{P_{i},C4^{\prime}_{i},P_{i+1},C4^{\prime}_{i+1}\}$ respectively, for each nucleotide $i$ along the chain. We observe that the dehedral angle distribution from RNA-FrameFlow closely recapitulates the distribution of naturally occuring RNA structures from the training set.
@@ 4.3 Generation quality across sequence lengths
We next investigate how sequence length affects the global realism of generated samples (measured by scTM). Figure 3 (Left) shows the performance of RNA-FrameFlow for different sequence lengths. We observe our model generates samples with high scTM for specific sequence lengths like 50, 60, 70, and 120 while generating poorer quality structures for other lengths. We believe the fluctuation of TM-scores may be due to certain lengths being over-represented in the training distribution. We can also partially attribute this to the inherent length bias of RhoFold; see Appendix B.2. With a better structure predictor, we expect an increase in valid samples that meet the 0.45 TM-score threshold.
We also analyze the novelty of our generated samples (measured by pdbTM) in Figure 3 (Middle). We are particularly interested in samples that lie in the right half with high scTM and low pdbTM, which means that the designs are highly likely to fold back into the sampled backbone but are structurally dissimilar to any RNAs in the training set. It is worth noting that our training set has high structural similarity among samples: running qTMclust on our training dataset revealed only 342 unique clusters from 5,319 samples, which indicates that the model does not encounter a diverse set of samples during training. This contributes to many generated samples from our model looking similar to samples from the training distribution. We include two such examples in Figure 3 (Right). Both generated RNAs yield relatively high pdbTM scores and look similar to their respective closest matching chain from the training set: a tRNA at length 70 and a 5S ribosomal RNA at length 120, respectively. We include comparative results on validity and novelty for MMDiff in Appendix D.1, finding that MMDiff does not generate any samples that pass the validity criteria.
Table 2: Local structural metrics. Earth Mover’s Distance for local structural measurements compared to ground truth measurements from RNAsolo. Our model ($N_{T}=50$) shows improved recapitulation of local structural descriptors compared to baselines.
@@
dist angles torsions
RNA-FrameFlow 0.17 0.11 2.36
@@ 4.4 Data preparation protocols
Due to the overrepresentation of RNA strands of certain lengths (mostly corresponding to tRNA or 5S ribosomal RNA) in our training set, our models generate close likenesses for those lengths that achieve high self-consistency but are not novel folds. To avoid this memorized recapitulation and promote increased diversity among samples, we sought to develop data preparation protocols to balance RNA folds across sequence lengths.
  • Structural clustering: We cluster our training set using qTMclust. When creating a training batch, we sample random clusters and from each, a random structure. This ensures a batch does not comprise solely of samples for a single sequence length or is dominated by over-represented folds. There are only 342 structural clusters for the 5,319 samples within sequence lengths 40-150, highlighting the lack of diversity in RNA structural data.
@@
We train identical models on these data splits for $120$K gradient steps, with evaluation results reported in Table 3 showing improved diversity and novelty in the generated samples, at the cost of reduced validity. For structural clustering, each batch comprises padded samples up to a maximum length of 150 from randomly selected structural clusters across sequence lengths. See Section D.2 for full results for the two alternate data preparation protocols.
@@ 5 Limitations and Discussions
Altogether, our experiments demonstrate that the $SE(3)$ flow matching framework is sufficiently expressive for learning the distribution of 3D RNA structure and generating realistic RNA backbones similar to well-represented RNA folds in the PDB. Select examples are shown in LABEL:fig:addition-samples. We have also identified notable limitations and avenues for future work, which we highlight below.
@@
Physical violations. While well-trained models usually generate realistic RNA backbones, we do observe some physical violations: generated backbones sometimes have chains that are either too close by or directly clash with one another, are highly coiled, have excessive loops and unrealistically intertwined helices, or have chain breaks. We highlight these limitations in Figure 5. RNA tertiary structure folding is driven by base pairing and base stacking (Vicens and Kieft, 2022) which influence the formation of helices, loops, and other tertiary motifs. Base pairing refers to nucleotides along adjacent chains forming hydrogen bonds, while base stacking involves interactions between rings of adjacent nucleotide bases along a chain. To our knowledge, all current deep learning models operate on individual nucleotides, only implicitly learning base pairing and stacking. Developing explicit representations of these interactions as part of the architecture may further minimize physical violations and provide stronger inductive biases to learn complex tertiary RNA motifs.
@@
Generalisation and novelty. We observed that the best designs from our models (as measured by scTM score) are sampled at lengths 70-80 and 120-130, and often have closely matching structures in the PDB (high TM-scores). This suggests that models can recapitulate well-represented RNA folds in their training distribution (e.g., both tRNAs at length 70-90 and small 5S ribosomal RNAs at length 120 are very frequent). However, self-consistency metrics were relatively poorer for less frequent lengths, suggesting that models are currently not designing novel folds.
@@
We would also like to note that the models we use for structure prediction and inverse folding may be similarly biased to perform well for certain sequence lengths, leading to the overall pipeline being reliable for commonly occurring lengths and unreliable for less frequent ones (see Appendix B.2 for an analysis on RhoFold). We evaluated preliminary strategies for structural clustering and cropping augmentations during training, which improved the novelty of designed structures but led to fewer designs passing the validity filter. Overall, the relative scarcity of RNA structural data compared to proteins necessitates greater care in preparing data pipelines for scaling up training and/or incorporating inductive biases into generative models, which we hope to continue exploring.
@@ 6 Conclusion
We introduce RNA-FrameFlow, a generative model for 3D RNA backbone design. Our evaluations show that our model can design locally realistic and moderately novel backbones of length 40 – 150 nucleotides. We achieve a validity score of 41.0% and relatively strong diversity and novelty scores compared to diffusion model baselines and ablated variants. While generative models can successfully recapitulate well-represented RNA folds (e.g., tRNAs, small rRNAs), the lack of diversity in the training data may hinder broad generalization at present. We are actively exploring improved data preparation strategies combined with inductive biases that explicitly incorporate physical interactions that drive RNA structure. We hope RNA-FrameFlow and the associated evaluation framework can serve as foundations for the community to explore 3D RNA design, towards developing conditional generative models for real-world design scenarios.
Acknowledgements
We would like to thank Jason Yim and Emile Mathieu for helpful comments and discussions. CKJ was supported by the A*STAR Singapore National Science Scholarship (PhD). AM was supported by a U.S. NSF grant (DBI2308699) and two U.S. NIH grants (R01GM093123 and R01GM146340). SVM was supported by the UKRI Centre for Doctoral Training in Application of Artificial Intelligence to the study of Environmental Risks (EP/S022961/1). This research was partially supported by a Cambridge Dawn Supercomputer Pioneer Project compute grant.
Figure 5: Physical violations in generated samples. (A) Steric clashes between generated chains (highlighted in yellow). (B) Chain breaks and stray strands (highlighted in yellow). (C)-(E) Excessive loops and intertwined helices.
Appendix A Related Work
Here, we summarize recent developments in deep learning for 3D RNA modeling and design. Recent end-to-end RNA structure prediction tools include RhoFold (Shen et al., 2022), RoseTTAFold2NA (Baek et al., 2022), DRFold (Li et al., 2023b), and AlphaFold3 (Abramson et al., 2024), each with varying performance that is yet to match the current state-of-the-art for proteins. Other approaches use GNNs as ranking functions (Townshend et al., 2021) together with sampling algorithms (Boniecki et al., 2016, Watkins et al., 2020). However, structure prediction tools are not directly capable of designing new structures, which this work aims to address by adapting an $SE(3)$ flow matching framework for proteins (Yim et al., 2023a). MMDiff (Morehead et al., 2023), a diffusion model for protein-nucleic acid complex generation, can also sample RNA-only structures in principle. Our evaluation shows that our flow matching model significantly outperforms both the original and RNA-only versions of MMDiff that we re-trained for fair comparison.
Joshi et al. (2023) introduce gRNAde, a GNN-based encoder-decoder for 3D RNA inverse folding, a closely related task of designing new sequences conditioned on backbone structures. Tan et al. (2023) and Shulgina et al. (2024) have also developed GNNs for 3D RNA inverse folding. We use gRNAde (Joshi et al., 2023) followed by RhoFold (Shen et al., 2022) in our evaluation pipeline to forward fold designed backbones and measure structural self-consistency.
Independently and concurrent to our work, Nori and Jin (2024) propose RNAFlow, an SE(3) flow matching model to co-design RNA sequence and structure conditioned on protein partners. At each denoising step, RNAFlow uses a protein-conditioned variant of gRNAde (Joshi et al., 2023) to inverse fold noised structures, followed by RoseTTAFold2NA (Baek et al., 2022) to predict the structure of the designed sequence. The performance of RNAFlow is upper-bounded by RoseTTAFold2NA as a pre-trained structure generator, which is kept frozen and not developed for designed RNAs which do not have co-evolutionary MSA information. Our work tackles de novo 3D RNA backbone generation, an orthogonal design task of sampling RNA backbone structures. We train RNA structure generation models from scratch, akin to recent developments in protein design (Yim et al., 2023b, a, Bose et al., 2023, Lin and AlQuraishi, 2023). Backbone generation followed by inverse folding has shown experimental success in designing functional proteins (Dauparas et al., 2022, Watson et al., 2023, Ingraham et al., 2022), as the framework is flexible for including specific structural motifs and sequence constraints.
Appendix B Additional Experimental Details
B.1 Hyperparameters for denoiser
Table 4: Hyperparameters for denoiser model.
@@ B.1 Hyperparameters for denoiser
Schedule Translations (training) linear
Rotations (training) linear
Translations (sampling) linear
Rotations (sampling) exponential
Number of denoising steps $N_{T}$ $[10,\mathbf{50},100,500]$
B.2 RhoFold Length Bias

### v3

@@ 4.1 Global evaluation of generated RNA backbones
We begin by analyzing RNA-FrameFlow’s samples using the aforementioned evaluation metrics. For validity, we report percentage of samples with scTM $\geq 0.45$; for diversity, we report the ratio of unique structural clusters to total valid samples; and for novelty, we report the highest average pdbTM to a match from the PDB. For each sequence length between 40 and 150, at intervals of 10, we generate 50 backbones. Table 1 reports these metrics across different variants for the number of denoising steps $N_{T}$. The average scTM and scRMSD of valid samples are $0.641\pm 0.161$ and $2.298\pm 0.892$ respectively. We compare our model to MMDiff (Morehead et al., 2023), a protein-RNA-DNA complex co-design diffusion model. As the original best-performing version of MMDiff was trained on shorted RNA sequences, we retrain it on our training set. We also inverse-fold MMDiff’s backbones using gRNAde.
We identify $N_{T}=50$ as the best-performing model that balances validity, diversity, and novelty; furthermore, it takes 4.74 seconds (averaged over 5 runs) to sample a backbone of length 100, as opposed to 27.3 seconds for MMDiff with 100 diffusion steps. We note that increasing $N_{T}$ does not improve validity despite allowing the model to perform more updates to atomic coordinate placements. Our model also outperforms MMDiff. On manual inspection, samples from MMDiff had significant chain breaks and disconnected floating strands; see Appendix C.1.
Table 1: Unconditional RNA backbone generation. We evaluate the performance of RNA-FrameFlow for multiple values for denoising steps $N_{T}$. The best-performing model uses $N_{T}=50$ steps, taking 4.74s to sample a backbone of length 100. We also provide the average self-consistency TM-score and RMSD value for all valid samples. We green-highlight the best result per column.
@@ 4.2 Local evaluation with structural measurements
For our best-performing model using diffusion timesteps $N_{T}=50$, we plot histograms of bond distance, bond angles, and dihedral angles in Figure 4 (Subplots 1-3). We include the Earth Mover’s distance (EMD) between measurements from the training and generated distributions as an indicator of local realism (using 30 bins for each quantity). An ideal generative model will score an EMD close to 0.0 (i.e. consistent with the training set comprising naturally occurring RNA). In Table 3, we observe EMD values from our best-performing model’s backbones being significantly closer to 0.0 compared to MMDiff. We include histograms of local structural descriptors for MMDiff in Appendix C.1.
We also show RNA Ramachandran angle plots for generated samples and the training distribution in Figure 4 (Subplot 4). Keating et al. (2011) introduced $\eta-\theta$ plots, similar to Ramachandran angle plots for proteins, that track the separate dihedral angles formed by $\{C4^{\prime}_{i},P_{i+1},C4^{\prime}_{i+1},P_{i+2}\}$ and $\{P_{i},C4^{\prime}_{i},P_{i+1},C4^{\prime}_{i+1}\}$ respectively, for each nucleotide $i$ along the chain. We observe that the dihedral angle distribution from RNA-FrameFlow closely recapitulates the angular distribution from naturally occurring RNA structures in the training set.
@@ 4.3 Generation quality across sequence lengths
We next investigate how sequence length affects the global realism of generated samples (measured by scTM). Figure 3 (Left) shows the performance of RNA-FrameFlow for different sequence lengths. We observe our model generates samples with high scTM for specific sequence lengths like 50, 60, 70, and 120 while generating poorer quality structures for other lengths. We believe the overrepresentation of certain lengths in the training distribution causes the fluctuation of TM-scores. We can also partially attribute this to the inherent length bias of RhoFold; see Appendix A.3. With better structure predictors, we expect more samples to be valid. We provide additional local evaluations of angular distributions in Appendix C.3.
We also analyze the novelty of our generated samples (measured by pdbTM) in Figure 3 (Middle). We are particularly interested in samples that lie in the right half with high scTM and low pdbTM, which means that the designs are highly likely to fold back into the sampled backbone but are structurally dissimilar to any RNAs in the training set. It is worth noting that our training set has high structural similarity among samples: running qTMclust on our training dataset revealed only 342 unique clusters from 5,319 samples, which indicates that the model does not encounter a diverse set of samples during training. This contributes to many generated samples from our model looking similar to samples from the training distribution. We include two such examples in Figure 3 (Right). Both generated RNAs yield relatively high pdbTM scores and look similar to their respective closest matching chain from the training set: a tRNA at length 70 and a 5S ribosomal RNA at length 120, respectively. We include comparative results on validity and novelty for MMDiff in Appendix C.1, finding that MMDiff does not generate any samples that pass the validity criteria.
Table 2: Local structural metrics. Earth Mover’s Distance for local structural measurements compared to ground truth measurements from RNAsolo. We also include EMD for a 50/50 train split as a sanity check. Our model shows improved recapitulation of local structural descriptors compared to baselines.
@@
distance angles torsions
50/50 training dist. $6.25\times 10^{-2}$ $8.97\times 10^{-4}$ $7.24\times 10^{-5}$
RNA-FrameFlow ($N_{T}=50$) 0.17 0.11 2.36
@@ 4.4 Data preparation protocols
Due to the overrepresentation of RNA strands of certain lengths (mostly corresponding to tRNA or 5S ribosomal RNA) in our training set, our models generate close likenesses for those lengths that achieve high self-consistency but are not novel folds. To avoid this memorized recapitulation and promote increased diversity among samples, we sought to develop data preparation protocols to balance RNA folds across sequence lengths. We identically train RNA-FrameFlow on these data splits for $120$K gradient steps, with results reported in Table 3.
  • Structural clustering: We cluster our training set using qTMclust. When creating a training batch, we sample random clusters, and from each cluster, random structures. This ensures batches do not solely contain samples for a single sequence length or are dominated by over-represented folds. There are only 342 structural clusters for the 5,319 samples within sequence lengths 40-150, highlighting the lack of diversity in RNA structural data. Each batch comprises padded samples up to a maximum length of 150 from randomly selected clusters across sequence lengths.
@@
We observe improved diversity and novelty at the cost of reduced validity. Randomly cropping may introduce subsequences that fold into significantly different structures than the substructure extracted from the original RNA; these subsequences may even unfold in real life. As a result, the augmented dataset may contain folds that are unstable or implausible. The structure prediction and inverse-folding models may not have encountered these folds loosely recapitulated by our model, resulting in poor validity. We are actively developing principled cropping methods that capture unique, realistic folds. We include additional results on these data preparation protocols in Appendix C.2.
Figure 5: Physical violations in generated samples. (A) Inter-chain clashes (highlighted yellow). (B) Chain breaks and stray strands (highlighted yellow). (C)-(E) Excessive loops and helices.
@@ 5 Limitations and Discussions
Altogether, our experiments demonstrate that the $SE(3)$ flow matching framework is sufficiently expressive for learning the distribution of 3D RNA structure and generating realistic RNA backbones similar to well-represented RNA folds in the PDB. Select examples are shown in Figure LABEL:fig:addition-samples. We have also identified notable limitations and avenues for future work, which we highlight below.
@@
Physical violations. While well-trained models usually generate realistic RNA backbones, we do observe some physical violations: generated backbones sometimes have chains that are either too close by or directly clash with one another, are highly coiled, have excessive loops and unrealistically intertwined helices, or have chain breaks. We highlight these limitations in Figure 5. RNA tertiary structure folding is driven by base pairing and base stacking which influence the formation of helices, loops, and other tertiary motifs (Vicens & Kieft, 2022) . Base pairing refers to nucleotides along adjacent chains forming hydrogen bonds, while base stacking involves interactions between rings of adjacent nucleotide bases along a chain. To our knowledge, all current deep learning models operate on individual nucleotides, only implicitly learning base pairing and stacking. Developing explicit representations of these interactions as part of the architecture may further minimize physical violations and provide stronger inductive biases to learn complex tertiary RNA motifs. We analyze steric clashes in our generated backbones in Appendix C.4.
@@
Generalization and novelty. We observed that the best designs from our models (as measured by scTM score) are sampled at lengths 70-80 and 120-130, and often have closely matching structures in the PDB (high TM-scores). This suggests that models can recapitulate well-represented RNA folds in their training distribution (e.g., both tRNAs at length 70-90 and small 5S ribosomal RNAs at length 120 are very frequent). However, self-consistency metrics were relatively poorer for less frequent lengths, suggesting that models are currently not designing novel folds.
@@
We would also like to note that the models we use for structure prediction and inverse folding may be similarly biased to perform well for certain sequence lengths, leading to the overall pipeline being reliable for commonly occurring lengths and unreliable for less frequent ones (see Appendix A.3 for an analysis on RhoFold). We evaluated preliminary strategies for structural clustering and cropping augmentations during training, which improved the novelty of designed structures but led to fewer designs passing the validity filter. Overall, the relative scarcity of RNA structural data compared to proteins necessitates greater care in preparing data pipelines for scaling up training and incorporating inductive biases into generative models.
6 Related Work
Recent end-to-end RNA structure prediction tools include RhoFold (Shen et al., 2022), RoseTTAFold2NA (Baek et al., 2022a), DRFold (Li et al., 2023a), and AlphaFold3 (Abramson et al., 2024), each with varying performance that is yet to match the current state-of-the-art for proteins. Other approaches use GNNs as ranking functions (Townshend et al., 2021) together with sampling algorithms (Boniecki et al., 2016; Watkins et al., 2020). However, structure prediction tools are not directly capable of designing new structures, which this work aims to address by adapting an $SE(3)$ flow matching framework for proteins (Yim et al., 2023a). MMDiff (Morehead et al., 2023), a diffusion model for protein-nucleic acid complex generation, can also sample RNA-only structures in principle. Our evaluation shows that our flow matching model significantly outperforms both the original and RNA-only versions of MMDiff that we re-trained for fair comparison.
Joshi et al. (2025) introduce gRNAde, a GNN-based encoder-decoder for 3D RNA inverse folding, a closely related task of designing new sequences conditioned on backbone structures. Tan et al. (2023) and Shulgina et al. (2024) have also developed GNNs for 3D RNA inverse folding. We use gRNAde (Joshi et al., 2025) followed by RhoFold (Shen et al., 2022) in our evaluation pipeline to forward fold designed backbones and measure structural self-consistency.
Independently and concurrent to our work, Nori & Jin (2024) propose RNAFlow, an SE(3) flow matching model to co-design RNA sequence and structure conditioned on protein partners. At each denoising step, RNAFlow uses a protein-conditioned variant of gRNAde (Joshi et al., 2025) to inverse fold noised structures, followed by RoseTTAFold2NA (Baek et al., 2022a) to predict the structure of the designed sequence. The performance of RNAFlow is upper-bounded by RoseTTAFold2NA as a pre-trained structure generator, which is kept frozen and not developed for designed RNAs which do not have co-evolutionary MSA information. Our work tackles de novo 3D RNA backbone generation, an orthogonal design task of sampling RNA backbone structures. We train RNA structure generation models from scratch, akin to recent developments in protein design (Yim et al., 2023b; a; Bose et al., 2023; Lin & AlQuraishi, 2023). Backbone generation followed by inverse folding has shown experimental success in designing functional proteins (Dauparas et al., 2022; Watson et al., 2023; Ingraham et al., 2022), as the framework is flexible for including specific structural motifs and sequence constraints.
@@ 7 Conclusion
We introduce RNA-FrameFlow, a generative model for 3D RNA backbone design. Our evaluations show that our model can design locally realistic and moderately novel backbones of length 40 – 150 nucleotides. We achieve a validity score of 41.0% and relatively strong diversity and novelty scores compared to diffusion model baselines and ablated variants. While generative models can successfully recapitulate well-represented RNA folds (e.g., tRNAs, small rRNAs), the lack of diversity in the training data may hinder broad generalization at present. Directions for future research include exploring improved data preparation strategies combined with inductive biases that explicitly incorporate physical interactions that drive RNA structure, including base pairing and base stacking. We hope RNA-FrameFlow and the associated evaluation framework can serve as foundations for the community to explore 3D RNA design, towards developing conditional generative models for real-world design scenarios.
Appendix A Additional Experimental Details
A.1 Denoiser Hyperparameters
@@
Schedule Translations (training / sampling) linear / linear
Rotations (training / sampling) linear / exponential
Number of denoising steps $N_{T}$ 50
Table 4: Hyperparameters for best performing denoiser model.
A.2 Upper bound performance of the Self-consistency Pipeline
Our self-consistency pipeline to compute validity involves inverse and forward folding using gRNAde (Joshi et al., 2025) and RhoFold (Shen et al., 2022). Placing upper bounds on the performance of our RNA backbone design pipeline offers insights into areas of improvement using available open-source tools.
To quantify the total error accumulated in our self-consistency pipeline, and its impact on downstream validity, we study the extent to which gRNAde and RhoFold can retrieve the ground truth sequences and structures from the RNAsolo training set. To assess RhoFold’s structure prediction performance, we take all sequences of length 40 – 150 from RNAsolo, forward-fold (FF) them using RhoFold, and compute self-consistency metrics (TM-score, RMSD) by comparing them with the sequences’ associated 3D folds. To assess gRNAde’s sequence recovery performance, we inverse-fold (IF) 3D backbones from RNAsolo through gRNAde to get 16 likely sequences and pass them to RhoFold for forward-folding.
As shown in the table below, the average self-consistency of the gRNAde-RhoFold pipeline with RNAsolo ground truth backbone structures is $43.7\%$, close to RNA-FrameFlow’s validity of $41.0\%$. This shows us that the generated backbones from RNA-FrameFlow closely retain the validity of RNAsolo backbones and corresponding sequences from gRNAde. In Figure 11, we also show the self-consistency TM-scores per length bins.
Pipeline Self-consistency (%) $\uparrow$ Avg scTM $\uparrow$ Avg scRMSD $\downarrow$
RNAsolo + FF only 55.1 0.690 2.804
RNAsolo + IF + FF 43.7 0.663 3.085
RNA-FrameFlow + IF + FF (ours) 41.0 0.641 2.298
Figure 11: Self-consistency scores on RNAsolo samples by sequence length. We observe that generated backbones from RNA-FrameFlow retain the self-consistency of gRNAde-predicted sequences.
A.3 RhoFold Length Bias